OpenStudy (anonymous):

Cos(theta)[sec(theta)-cos(theta)]=sin^2(theta)

rishavraj (rishavraj):

\[a(b - c) = ab - ac \] \[\cos \theta \times \sec \theta = 1 \]

OpenStudy (anonymous):

this ones easy

OpenStudy (anonymous):

just put sec theta = to 1/cos theta then simply multiply

OpenStudy (anonymous):

I don't understand your way lol

OpenStudy (anonymous):

it will become 1 - cos^2 theta which is equal to sin^2 theta

OpenStudy (anonymous):